OpenStudy (anonymous):

9over1300 convert to mixed number

OpenStudy (whpalmer4):

\[\frac{9}{1300}\]or\[\frac{1300}{9}\]The first one doesn't make any sense as a mixed number, but that's what "9 over 1300" means

OpenStudy (whpalmer4):

Say what?! \[\frac{1300}{9}\]\[1300/9 = 144.444\]so the whole number portion is 144. Multiply 144*9 and subtract the result from 1300 to get the numerator of fractional part. For example, 120/7 as a mixed number: \[120/7 = 17.142857...\]so the whole number is 17\[120-17*7 = 1\]so the fraction's numerator is 1, and denominator is 7\[\frac{120}{7} = 17\frac{1}{7}\]